You have just been hired as a new management trainee by Earrings Unlimited, a distributor of earrings to various retail outlets located in shopping malls across the country. In the past, the company has done very little in the way of budgeting and at certain times of the year has experienced a shortage of cash. Since you are well trained in budgeting, you have decided to prepare a master budget for the upcoming second quarter. To this end, you have worked with accounting and other areas to gather the information assembled below.

January (actual) 22,000 June (budget) 52,000
February (actual) 28,000 July (budget) 32,000
March (actual) 42,000 August (budget) 30,000
April (budget) 67,000 September (budget) 27,000
May (budget) 102,000

The company sells many styles of earrings, but all are sold for the same price—$16 per pair. Actual sales of earrings for the last three months and budgeted sales for the next six months follow (in pairs of earrings):

The concentration of sales before and during May is due to Mother’s Day. Sufficient inventory should be on hand at the end of each month to supply 40% of the earrings sold in the following month.

Suppliers are paid $5.00 for a pair of earrings. One-half of a month’s purchases is paid for in the month of purchase; the other half is paid for in the following month. All sales are on credit. Only 20% of a month’s sales are collected in the month of sale. An additional 70% is collected in the following month, and the remaining 10% is collected in the second month following sale. Bad debts have been negligible.

Monthly operating expenses for the company are given below:

Variable:
Sales commissions 4 % of sales
Fixed:
Advertising $ 300,000
Rent $ 28,000
Salaries $ 126,000
Utilities $ 12,000
Insurance $ 4,000
Depreciation $ 24,000

Insurance is paid on an annual basis, in November of each year.

The company plans to purchase $21,000 in new equipment during May and $50,000 in new equipment during June; both purchases will be for cash. The company declares dividends of $22,500 each quarter, payable in the first month of the following quarter.

The company’s balance sheet as of March 31 is given below:

Assets
Cash $ 84,000
Accounts receivable ($44,800 February sales; $537,600 March sales) 582,400
Inventory 134,000
Prepaid insurance 26,000
Property and equipment (net) 1,050,000
Total assets $ 1,876,400
Liabilities and Stockholders’ Equity
Accounts payable $ 110,000
Dividends payable 22,500
Common stock 1,000,000
Retained earnings 743,900
Total liabilities and stockholders’ equity $ 1,876,400

The company maintains a minimum cash balance of $60,000. All borrowing is done at the beginning of a month; any repayments are made at the end of a month.

The company has an agreement with a bank that allows the company to borrow in increments of $1,000 at the beginning of each month. The interest rate on these loans is 1% per month and for simplicity we will assume that interest is not compounded. At the end of the quarter, the company would pay the bank all of the accumulated interest on the loan and as much of the loan as possible (in increments of $1,000), while still retaining at least $60,000 in cash.

Required:

Prepare a master budget for the three-month period ending June 30. Include the following detailed schedules:

1. a. A sales budget, by month and in total.

    b. A schedule of expected cash collections, by month and in total.

    c. A merchandise purchases budget in units and in dollars. Show the budget by month and in total.

    d. A schedule of expected cash disbursements for merchandise purchases, by month and in total.

2. A cash budget. Show the budget by month and in total. Determine any borrowing that would be needed to maintain the minimum cash balance of $60,000.

3. A budgeted income statement for the three-month period ending June 30. Use the contribution approach.

4. A budgeted balance sheet as of June 30.
Answer #1

A C D E F G HI 1 Sales Budget April 67000 May 102000 3 4 5 June 52000 16 832000 Units sold S.P Sales I Quarter 221000 16 3536DE Cash Budget April May 84,000 60,720 7,29,600 11,44,000 8.13,600 12.04,720 June Quarter 2,44,940 84,000 14,16,000 32,89,600А 58 Earrings Unlimited 59 Income Statement 60 For the Quarter Ended June 30 61 35,36,000 11,05,000 1.41.440 (134000 + 103500C D 7,02,930 8,28,800 (1632000 10% + 832000 * 80% ) 64,000 Closing inventory = 12800 * 5 = 64000 14,000 (26000 - 12000) 16,09Sales Budget May Quarter 67000 102000 52000 =B3+C3+D3 3 4 5 Units sold S.P Sales 16 16 =B3 B4 =C3*C4 =D3 D4 =E3 E4 Schedule oCash Budget June =C56 April 84000 =B14 =B37+B38 May =B56 =C14 =C37+C38 Quarter =B37 =E14 =E37+E38 = 14 =D37+D38 37 Cash BalanD 58 Earrings Unlimited 59 Income Statement 60 For the Quarter Ended June 30 61 -E5 1105000 =E45 =B64+B65 =C62-C65 ( 134000 +=D56 828800 64000 14000 =C81+C82+C83+C84 A 77 78 Balance Sheet 79 Assets 80 Current Assets: 81 Cash 82 Accounts Receivable 83